OpenStudy (freemap):

What is the distance between E(5, 1) and F(7, 10)? Round to the nearest tenth.

OpenStudy (freemap):

9 is my answer

OpenStudy (michele_laino):

hint: we can write this: \[\begin{gathered} d = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} = \hfill \\ \hfill \\ = \sqrt {{{\left( {5 - 7} \right)}^2} + {{\left( {1 - 10} \right)}^2}} = \hfill \\ \hfill \\ = \sqrt {{{\left( { - 2} \right)}^2} + {{\left( { - 9} \right)}^2}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (freemap):

4+81=85 9.219

OpenStudy (michele_laino):

that's right! Now, please round such value as requested

OpenStudy (freemap):

9.3

OpenStudy (freemap):

no 9.2

OpenStudy (michele_laino):

correct!

OpenStudy (freemap):

Thanks

OpenStudy (michele_laino):

:)